OpenStudy (anonymous):

MEDAL & FAN. Solve for x: 2 over 5 (x - 2) = 4x. 2 over 9 9 negative 2 over 9 negative 9 over 2

OpenStudy (anonymous):

I don't think if this is the answer.. \[\frac{ 2 }{ 5 }(x-2) = 4x\] \[0.4(x-2) = 4x\] \[0.4x - 0.8 = 4x\] \[-0.8 = -0.4x + 4x\] \[-0.8 = 3.6x\] \[\frac{ -0.8 }{ 3.6 } = \frac{ 3.6x }{ 3.6 }\] \[0.22 = x\] \[2/9 = x\] Now i use to divide first the 2/5 and get the answer on it,next i multiply 0.4.. See on the top my solution..